OpenStudy (anonymous):

Sarah is the president of the drama club and wants to publish a newspaper. Expenses are $.90 for printing and mailing each copy, plus $600 total for research and writing. The price of the newsletter is $1.50 per copy. How many copies of the newsletter must the club sell to break even? a. To break even, the drama club must sell 1000 copies. b. To break even, the drama club must sell 100 copies. c. To break even, the drama club must sell 250 copies. d. To break even, the drama club must sell 444 copies.

OpenStudy (anonymous):

.90x + 600 = 1.50x -- subtract .90 from both sides 600 = 1.50x - .90x 600 = .60x --- now divide by .60 600/.60 = x 1000 = x check.. .90(1000) + 600 = 1.50(1000) 900 + 600 = 1500 1500 = 1500 (correct) To break even, the drama club must sell 1000 copies.

OpenStudy (anonymous):

thank you so much kelliegirl33

OpenStudy (anonymous):

no problem :)